eviechocolate:

A student expanded an expression, as shown. Is the student's work correct? Explain why or why not. −6 ( 4x − 2/13)−6(4x) + 6 ( − 2/13 ) −24x − 12/13

AZ:

You need to know the distributive property \( a(b-c) = ab - ac\) Now take a look at the work that the student did Pay extra attention to the signs and the plus or minuses in the expression We have -6 and we're distributing that inside This is also an important concept that you should know: positive number * positive number = positive number negative number * negative number = positive number positive number * negative number = negative number negative number * positive number = negative number

AZ:

re-writing the work that the student did for clarity \(−6 \left(4x − \dfrac{2}{13}\right) \) \(−6(4x) + 6\left( −\dfrac{2}{13}\right) \) \( −24x − \dfrac{12}{13}\)

AZ:

So let me add colors for you \(\large \color{orange}{a}(b\color{red}{−}c) = \color{orange}{a}b \color{red}{−} \color{orange}{a}c\) So in our question, the 'a' is -6 But do you see how that minus sign stays? You can later simplify depending on the value of 'c' using the properties I mentioned above

AZ:

So did the student make a mistake? And if so, where?